# On Robust hypothesis testing with respect to Hellinger distance

### v2

@@ 5 Some results on testing with respect to symmetric $\chi^{2}$ distance
$\mathbb{E}[T(X)]$ $\geq(2-\sqrt{2})H^{2}(p,p_{2})-(2+\sqrt{2})H^{2}(p,p_{1})$
$\geq\frac{2-\sqrt{2}}{4}\chi^{2}(p,p_{2})-\frac{2+\sqrt{2}}{2}\chi^{2}(p,p_{1})$
Thus, $\mathbb{E}[T(X)]\geq 0$ if $\chi^{2}(p,p_{2})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{1})$. Likewise, we can show that $\mathbb{E}[T(X)]\leq 0$ if $\chi^{2}(p,p_{1})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{2})$. We can get a lower bound of $3$ for robust symmetric $\chi^{2}$ testing. The construction of the lower bound is exactly the same (our choice of parameters is $b=1,a_{2}=1,a_{1}\rightarrow\infty$. We only need to show the following. For any distribution $p$ in the perturbed family around $p_{1}$, we have
$\chi^{2}(p,p_{1})$ $=\frac{(a_{1}-0)^{2}}{a_{1}}\left(\frac{1}{2N_{m}}\right)\left(\frac{1}{a_{1}}N_{m}\right)+\frac{(2-1)^{2}}{2+1}\left(\frac{1}{2N_{m}}\right)N_{m}$
@@
$\chi^{2}(p,p_{2})$ $=\frac{(2-0)^{2}}{2}\frac{1}{2N_{m}}\left(1-\frac{1}{a_{1}}\right)N_{m}+\frac{(a_{1}-2)^{2}}{a_{1}+2}\frac{1}{2N_{m}}\frac{1}{a_{1}}N_{m}+\frac{(1-0)^{2}}{1}\frac{1}{2N_{m}}N_{m}$
@@
$\lim_{a_{1}\rightarrow\infty}\frac{\chi^{2}(p,q_{2})}{\chi^{2}(p,q_{1})}$ $=\frac{2+1+1}{\frac{4}{3}}$
@@
Likewise, for any distribution $p$ in the perturbed family around $p_{2}$. The rest of the proof is similar to the Hellinger case.
6 Alternate Test for the Composite Testing Problem
Consider the following composite hypothesis test. Let $p_{1}$ and $p_{2}$ be our hypotheses. The true distribution $p$ could lie in a Hellinger squared ball of radius $r$ around $p_{1}$ or $p_{2}$. As before, $p_{1}$, $p_{2}$, and $p$ are on some common domain $\mathcal{X}$.
$H_{0}$ $:H^{2}(p,p_{1})\leq r$
$H_{1}$ $:H^{2}(p,p_{2})\leq r.$
Naturally, the radius $r$ should be such that the two balls do not intersect. Let $r^{*}$ be the radius when the two balls intersect. We want a test which distinguishes between the two hypotheses as long as r<rr<r*. Observe that $H_{0}$ and $H_{1}$ are convex sets. Hence, it is known that the approach of working with likelihood ratio tests with respect to the closest pair works [11, Chapter 7]. We show that a simple modification of Baraud’s test (1) can distinguish between $H_{0}$ and $H_{1}$ as long as $r<r^{*}$. We now characterize $r^{*}$. Recall that the Hellinger distance is related to the Bhattacharya distance in the following manner:
$H^{2}(p_{1},p_{2})$ $=1-B(p_{1},p_{2})$
$=1-\sum_{x\in\mathcal{X}}\sqrt{p_{1}(x)p_{2}(x)}.$
Since $\sqrt{p_{1}}$ and $\sqrt{p_{2}}$ reside on a unit $\ell_{2}$ ball, we denote $\cos\theta:=B(p_{1},p_{2})$. A geodesic (shortest path) in the Hellinger distance from $p_{1}$ and $p_{2}$ is given by [11, Chapter 7]
$\sqrt{q_{\phi}}=\frac{\sin{(\theta-\phi)}\sqrt{p_{1}}+\sin{\phi}\sqrt{p_{2}}}{\sin{\theta}}.$
It can be verified that $q_{\phi}$ is a valid probability distribution using the fact that $\sin^{2}{(\theta-\phi)}+\sin^{2}{\phi}+2\sin{(\theta-\phi)}\sin{\phi}\cos{\theta}=\sin^{2}{\theta}$. Here, $\phi\in[0,\theta]$ with $q_{0}=p_{1}$ and $q_{\theta}=p_{2}$. Thus, we can define $u$ to be the "Hellinger midpoint" of $p_{1}$ and $p_{2}$, that is,
$\sqrt{u}$ $:=q_{\frac{\theta}{2}}$
$=\frac{\sin{\frac{\theta}{2}}}{\sin{\theta}}(\sqrt{p_{1}}+\sqrt{p_{2}}).$
@@ 6 Alternate Test for the Composite Testing Problem
$r^{*}$ $=H^{2}(p_{1},u)$
$=1-\sum_{x\in\mathcal{X}}\sqrt{p_{1}(x)}\sqrt{u(x)}$
@@
$T(X^{n}):=\sum_{i=1}^{n}\left(\sqrt{\frac{p_{1}(X_{i})}{u(X_{i})}}-\sqrt{\frac{p_{2}(X_{i})}{u(X_{i})}}\right)$
We now analyze the expected value of $T(X^{n})$ with respect to the unknown distribution $p$. By linearity of expectation, it suffices to analyze for $n=1$.
$E[T(X)]=\sum_{x\in\mathcal{X}}p(x)\left(\sqrt{\frac{p_{1}(x)}{u(x)}}-\sqrt{\frac{p_{2}(x)}{u(x)}}\right).$
We now add and subtract $2\sum\sqrt{p(x)p_{1}(x)}$, $2\sum\sqrt{p(x)p_{2}(x)}$ and $\sum\sqrt{p_{1}(x)u(x)}$ (the last term is also equal to $\sum\sqrt{p_{2}(x)u(x)}$ by our choice of $u$). Thus, we have
@@
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}+\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{1}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}-\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{2}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}+\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{1}(x)}-\sqrt{p_{2}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}+\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{1}(x)}+\sqrt{p_{2}(x)}-2\sqrt{p_{2}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$\overset{(a)}{=}2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}+\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{1}(x)}+\sqrt{p_{2}(x)}-2\sqrt{p_{2}(x)}}{\sqrt{p_{1}(x)}+\sqrt{p_{2}(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$\overset{(b)}{\geq}2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}-\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}\sum_{x\in\mathcal{X}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}-2\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}(1-\sum_{x\in\mathcal{X}}\sqrt{p(x)u(x)})$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{2}(x)}-2\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}\left(1-\sum_{x\in\mathcal{X}}\sqrt{p(x)}\frac{\sin{\frac{\theta}{2}}}{\sin{\theta}}(\sqrt{p_{1}(x)}+\sqrt{p_{2}(x)})\right)$
$=4\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}$
$=4\left(\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-\cos{\frac{\theta}{2}}\right)$
$(a)$ follows from the definition of $\sqrt{u}$, $(b)$ follows from the fact that $\frac{\sqrt{p_{2}(x)}}{\sqrt{p_{1}(x)}+\sqrt{p_{2}(x)}}\leq 1$. Thus, if $\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}>\cos{\frac{\theta}{2}}$, $E[T(X)]>0$. If $\sum\sqrt{pp_{2}}>\cos{\frac{\theta}{2}}$, $E[T(X)]<0$. Or, in other words, if $H^{2}(p,p_{1})<1-\cos{\frac{\theta}{2}}$, then $E[T(X)]>0$. If $H^{2}(p,p_{2})<1-\cos{\frac{\theta}{2}}$, then $E[T(X)]<0$. Thus, our test can distinguish between $H_{0}$ and $H_{1}$ as long as $r<r^{*}=1-\cos{\frac{\theta}{2}}$.
@@ 7 Discussion
8 Acknowledgements
We thank Ananda Theertha Suresh and Sivaraman Balakrishnan for introducing us to this problem and for their helpful discussions in the initial part of this project.

### v3

@@ 5 Lower bound on testing with respect to symmetric $\chi^{2}$ distance
$\mathbb{E}[T(X)]$ $\geq(2-\sqrt{2})H^{2}(p,p_{1})-(2+\sqrt{2})H^{2}(p,p_{0})$
$\geq\frac{2-\sqrt{2}}{4}\chi^{2}(p,p_{1})-\frac{2+\sqrt{2}}{2}\chi^{2}(p,p_{0})$
Thus, $\mathbb{E}[T(X)]\geq 0$ if $\chi^{2}(p,p_{1})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{0})$. Likewise, we can show that $\mathbb{E}[T(X)]\leq 0$ if $\chi^{2}(p,p_{0})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{1})$. We can get a lower bound of $3$ for robust symmetric $\chi^{2}$ testing with respect to the symmetric $\chi^{2}$ distance.
Theorem 3.
For every $\gamma<3$, there is a class $\mathcal{P}=\{p_{0},p_{1}\}$ which is not $\gamma$-robustly testable (with respect to the symmetric $\chi^{2}$ distance).
Proof.
The construction of the lower bound is exactly the same (our choice of parameters is $b=1,a_{2}=1,a_{1}\rightarrow\infty$. We only need to show the following. For any distribution $p$ in the perturbed family around $p_{0}$, we have
$\chi^{2}(p,p_{0})$ $=\frac{(a_{1}-0)^{2}}{a_{1}}\left(\frac{1}{2N_{m}}\right)\left(\frac{1}{a_{1}}N_{m}\right)+\frac{(2-1)^{2}}{2+1}\left(\frac{1}{2N_{m}}\right)N_{m}$
@@
$\chi^{2}(p,p_{1})$ $=\frac{(2-0)^{2}}{2}\frac{1}{2N_{m}}\left(1-\frac{1}{a_{1}}\right)N_{m}+\frac{(a_{1}-2)^{2}}{a_{1}+2}\frac{1}{2N_{m}}\frac{1}{a_{1}}N_{m}+\frac{(1-0)^{2}}{1}\frac{1}{2N_{m}}N_{m}$
@@
$\lim_{a_{1}\rightarrow\infty}\frac{\chi^{2}(p,p_{1})}{\chi^{2}(p,p_{0})}$ $=\frac{2+1+1}{\frac{4}{3}}$
@@
Likewise, for any distribution $p$ in the perturbed family around $p_{1}$. The rest of the proof is similar to the Hellinger case. ∎
6 Alternative Formulation of Robustness
In previous sections, we let the two hypotheses as $H_{0}:\gamma d(p,p_{0})<d(p,p_{1})$ and $H_{1}:d(p,p_{0})>\gamma d(p,p_{1})$. An alternative way to define robustness is by letting each hypothesis be a Hellinger ball around a fixed distribution. More concretely, let $p_{0}$ and $p_{1}$ be our fixed distributions. The true distribution $p$ could lie in a Hellinger squared ball of radius $r$ around $p_{0}$ or $p_{1}$. As before, $p_{0}$, $p_{1}$, and $p$ are on some common domain $\mathcal{X}$.
$H_{0}$ $:H^{2}(p,p_{0})\leq r$
$H_{1}$ $:H^{2}(p,p_{1})\leq r.$
Naturally, the radius $r$ should be such that the two balls do not intersect. Let $r^{*}$ be the radius when the two balls intersect. We want a test which distinguishes between the two hypotheses as long as r<rr<r*. Observe that $H_{0}$ and $H_{1}$ are convex sets. Hence, it is known that the approach of working with likelihood ratio tests with respect to the closest pair works [4, Chapter 7]. We show that a simple modification of Baraud’s test (1) can distinguish between $H_{0}$ and $H_{1}$ as long as $r<r^{*}$. We now characterize $r^{*}$. Recall that the Hellinger distance is related to the Bhattacharya distance in the following manner:
$H^{2}(p_{0},p_{1})$ $=1-B(p_{0},p_{1})$
$=1-\sum_{x\in\mathcal{X}}\sqrt{p_{0}(x)p_{1}(x)}.$
Since $\sqrt{p_{0}}$ and $\sqrt{p_{1}}$ reside on a unit $\ell_{2}$ ball, we denote $\cos\theta:=B(p_{0},p_{1})$. A geodesic (shortest path) in the Hellinger distance from $p_{0}$ and $p_{1}$ is given by [4, Chapter 7]
$\sqrt{q_{\phi}}=\frac{\sin{(\theta-\phi)}\sqrt{p_{0}}+\sin{\phi}\sqrt{p_{1}}}{\sin{\theta}}.$
It can be verified that $q_{\phi}$ is a valid probability distribution using the fact that $\sin^{2}{(\theta-\phi)}+\sin^{2}{\phi}+2\sin{(\theta-\phi)}\sin{\phi}\cos{\theta}=\sin^{2}{\theta}$. Here, $\phi\in[0,\theta]$ with $q_{0}=p_{0}$ and $q_{\theta}=p_{1}$. Thus, we can define $u$ to be the "Hellinger midpoint" of $p_{0}$ and $p_{1}$, that is,
$\sqrt{u}$ $:=\sqrt{q_{\frac{\theta}{2}}}$
$=\frac{\sin{\frac{\theta}{2}}}{\sin{\theta}}(\sqrt{p_{0}}+\sqrt{p_{1}}).$
@@ 6 Alternative Formulation of Robustness
$r^{*}$ $=H^{2}(p_{0},u)$
$=1-\sum_{x\in\mathcal{X}}\sqrt{p_{0}(x)}\sqrt{u(x)}$
@@
$T(X^{m}):=\frac{1}{m}\sum_{i=1}^{m}\left(\sqrt{\frac{p_{0}(X_{i})}{u(X_{i})}}-\sqrt{\frac{p_{1}(X_{i})}{u(X_{i})}}\right)$ (2)
Theorem 4.
The test described in (2) can distinguish between $H_{0}$ and $H_{1}$ as long as $r<1-\cos\frac{\theta}{2}$ where $\theta=\cos^{-1}B(p_{0},p_{1})$.
Proof.
We analyze the expected value of $T(X^{m})$ with respect to the unknown distribution $p$. By linearity of expectation, it suffices to analyze for $m=1$.
$E[T(X)]=\sum_{x\in\mathcal{X}}p(x)\left(\sqrt{\frac{p_{0}(x)}{u(x)}}-\sqrt{\frac{p_{1}(x)}{u(x)}}\right).$
We now add and subtract $2\sum\sqrt{p(x)p_{0}(x)}$, $2\sum\sqrt{p(x)p_{1}(x)}$ and $\sum\sqrt{p_{0}(x)u(x)}$ (the last term is also equal to $\sum\sqrt{p_{1}(x)u(x)}$ by our choice of $u$). Thus, we have
@@
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}+\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{0}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}-\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{1}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}+\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{0}(x)}-\sqrt{p_{1}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}+\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{0}(x)}+\sqrt{p_{1}(x)}-2\sqrt{p_{1}(x)}}{\sqrt{u(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$\overset{(a)}{=}2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}+\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}\sum_{x\in\mathcal{X}}\frac{\sqrt{p_{0}(x)}+\sqrt{p_{1}(x)}-2\sqrt{p_{1}(x)}}{\sqrt{p_{0}(x)}+\sqrt{p_{1}(x)}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$\overset{(b)}{\geq}2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}\sum_{x\in\mathcal{X}}(\sqrt{p(x)}-\sqrt{u(x)})^{2}$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}(1-\sum_{x\in\mathcal{X}}\sqrt{p(x)u(x)})$
$=2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{1}(x)}-2\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}\left(1-\sum_{x\in\mathcal{X}}\sqrt{p(x)}\frac{\sin{\frac{\theta}{2}}}{\sin{\theta}}(\sqrt{p_{0}(x)}+\sqrt{p_{1}(x)})\right)$
$=4\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-2\frac{\sin{\theta}}{\sin{\frac{\theta}{2}}}$
$=4\left(\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}-\cos{\frac{\theta}{2}}\right)$
$(a)$ follows from the definition of $\sqrt{u}$, $(b)$ follows from the fact that $\frac{\sqrt{p_{1}(x)}}{\sqrt{p_{0}(x)}+\sqrt{p_{1}(x)}}\leq 1$. Thus, if $\sum_{x\in\mathcal{X}}\sqrt{p(x)p_{0}(x)}>\cos{\frac{\theta}{2}}$, $E[T(X)]>0$. If $\sum\sqrt{pp_{1}}>\cos{\frac{\theta}{2}}$, $E[T(X)]<0$. Or, in other words, if $H^{2}(p,p_{0})<1-\cos{\frac{\theta}{2}}$, then $E[T(X)]>0$. If $H^{2}(p,p_{1})<1-\cos{\frac{\theta}{2}}$, then $E[T(X)]<0$. Thus, our test can distinguish between $H_{0}$ and $H_{1}$ as long as $r<r^{*}=1-\cos{\frac{\theta}{2}}$. ∎
@@ 7 Discussion
8 Acknowledgments
We thank Ankit Pensia for bringing to our attention the following observation. Let $d(.,.)$ be any distance metric. Consider a different formulation of the robust testing problem. $H_{0}^{\prime}:d(p,p_{0})\leq\frac{d(p_{0},p_{1})}{C}$ vs $H_{1}^{\prime}:d(p,p_{1})\leq\frac{d(p_{0},p_{1})}{C}$. Birge (cite?) gives a test that works for any $C>2$. This implies a test for the problem $H_{0}:(C+1)d(p,p_{0})\leq d(p,p_{1})$ vs $H_{1}:d(p,p_{0})\geq(C+1)d(p,p_{1})$. Observe that
$(C+1)d(p,p_{0})$ $\leq d(p,p_{1})\leq d(p,p_{0})+d(p_{0},p_{1})$
$Cd(p,p_{0})$ $\leq d(p_{0},p_{1}).$
Thus, $H_{0}\implies H_{0}^{\prime}$. Likewise, $H_{1}\implies H_{1}^{\prime}$. Since Birge’s test works for any $C>2$, it also works for any $C+1>3$ for our problem. In particular, for the Hellinger distance, this bound is better than the one in [9] but not as good as the one in [1].